lan 5 15 The pdf of x is f(x) = 0.1, 3<x< 13. Find the 60th percentile. (the answer is an integer)
Answer #1

60th percentile means probability is 0.6

PX <r) = 0.6

13-1 = 0.6

X-3 = 10 * 0.6

x = 3 + 6 = 9

So, 9 is the answer.

Please comment if any doubt.